OpenStudy (anonymous):

Algebra 9(3x+9)=14-(x=3) help please

OpenStudy (anonymous):

Solving for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it \[9(3x+9)=14-x(-3)?\]

OpenStudy (anonymous):

-(x-3)*

OpenStudy (anonymous):

there is something wrong in your question!

OpenStudy (anonymous):

its (x+3)

OpenStudy (anonymous):

ok let me write it again 9(3x+9)=14-(x+3) solve for x

OpenStudy (anonymous):

x= -16/7 is the answer

OpenStudy (anonymous):

okay ??

OpenStudy (anonymous):

are u der @ wan?

OpenStudy (anonymous):

sorry thinker stepped away I have another problem 7x-7y=1 y=?

OpenStudy (anonymous):

ok....y=(1+7x)/7

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

so y =7

OpenStudy (anonymous):

no y=\[(1+7x)\div 7\]

OpenStudy (anonymous):

ok i think

OpenStudy (anonymous):

was ur equation 7x-7y=1 ??? and u need to find 'y' ???

OpenStudy (anonymous):

thinker what does it mean when they say subtract from both sides

OpenStudy (anonymous):

means, u subtract 'something' on left & right side of the '=' sign :) got it?

OpenStudy (anonymous):

and this 'something' may be any number

OpenStudy (anonymous):

ok..bye for now..all the best :)

OpenStudy (anonymous):

u got sumthing to say? be quick

OpenStudy (anonymous):

nothing..ok..bye then:)